OpenStudy (anonymous):

how do you transform 1-2cos^2A+(1-sin^2A)^2/cos^4 to (2-2cos^4A-2sin^2A+sin^4A)/cos4A

OpenStudy (anonymous):

Please Loser66 be my hero :D

OpenStudy (anonymous):

huh? I tried 1-2cos^2A+1-2sin^2A+sin^4A/cos^4 how ever from there I don't know what to do

OpenStudy (loser66):

let re write yours. \[\frac{1-2cos^2A+(1-sin^2A)^2}{cos^4A}\]right? is it yours?

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

\[(1-sin^2A) = cos^2A\\therefore~yours~will~become~ \frac{1-2cos^2A +cos^4A}{cos^4A}\]right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait...

OpenStudy (loser66):

yours second one is \[\frac{2-2cos^4A-2sin^2A+sin^4A}{cos^4A}\]my question is the denominator is cos^4 A or cos (4A)

OpenStudy (anonymous):

cos^4 A

OpenStudy (loser66):

nope, it is (1-cos^2A)^2

OpenStudy (anonymous):

let me re arrange them

OpenStudy (loser66):

but I cannot get your second one. hehehe... my turns to tan^4 A

OpenStudy (anonymous):

it is supposed to go to tan^4 A

OpenStudy (anonymous):

I'm just trying to make it larger

OpenStudy (loser66):

I got tan^4 A only. hehehe... sorry.

OpenStudy (anonymous):

but how?

OpenStudy (loser66):

\[\frac{1-2cos^2A +cos^4A}{cos^4A}= \frac{(1-sin^2A)^2}{cos^4A}=\frac{sin^4A}{cos^4A}=tan^4A\]

OpenStudy (loser66):

\[1-2cos^2A +cos^4A\]form the form of (a -2ab+b^2 ) =(a-b)^2 your a =1 your b =cos^2

OpenStudy (loser66):

I have to go.

OpenStudy (anonymous):

however how can you transform \[\frac{ 1-2 \cos ^{2}A+1-2\sin ^{2}A+\sin ^{4}A}{ \cos ^{4}A }\]

OpenStudy (loser66):

from numerator,\[1-2cos^2A+1-2sin^2A+sin^4A\\= 2-2(cos^2 A+sin^2A)+sin^4A\\= 2-2+sin^4A \\= sin^4A\] but I really not know why do you have to go on that way. Sorry for suddenly logged out, this morning I was late for class.